# FLOPS: Forward Learning with OPtimal Sampling

### v1

@@ C.4 Extended ablation
We provide a more extensive ablation study on ViT-Large following the same setting as the main text. Figure 5(a)-5(c) and Figure 5(d)-5(f) show the result of LR and SPSA estimator. Both the estimators can benefit from the allocation and the Gaussain allocator has better performance than the Bernoulli allocator. The gradient of deep layer like layer 1 is more difficult to estimate than the shallow layer and the corresponding variance accounts for most of the variance of the whole model. Comparing Figure 5(a)-5(c) to Figure 5(g)-5(i), estimation on the feed-forward network is easier than on the multi-head attention. In practice, it is recommended to add gradient clip to the multi-head attention layer.

### v3

@@ C.4 Extended ablation
We provide a more extensive ablation study on ViT-Large following the same setting as the main text. Figures 6(6(a))-(6(c)) and (6(d))-(6(f)) show the result of LR and SPSA estimator, respectively. Both the estimators can benefit from the allocation and the Gaussain allocator has better performance than the Bernoulli allocator. The gradient of deep layer like layer 1 is more difficult to estimate than the shallow layer and the corresponding variance accounts for most of the variance of the whole model. Comparing Figures 6(6(a))-(6(c)) to (6(g))-(6(i)), estimation on the feed-forward network is easier than on the multi-head attention. In practice, it is recommended to add gradient clipping to the multi-head attention layer.